# FERGI: Automatic Annotation of User Preferences for Text-to-Image Generation from Spontaneous Facial Expression Reaction

### 05 Dec 2023 v1

@@ 5.2.3 Baseline Scores
To compare the performance of our AU4 valence score, we use the CLIP score [55], Aesthetic score [60], BLIP score [40], ImageReward score [73], PickScore [32], and HPS v2 score [71] as baselines.
@@
$s_{\mathrm{ens}}=w_{\mathrm{IR}}s_{\mathrm{IR}}+w_{\mathrm{Pick}}s_{\mathrm{Pick}}+w_{\mathrm{HPSv2}}s_{\mathrm{HPSv2}}$
@@
To be consistent with the evaluation of AU4 valence score, a leave-one-participant-out (LOPO) procedure was applied in the fitting of the three weights $w_{\mathrm{IR}}$, $w_{\mathrm{Pick}}$, and $w_{\mathrm{HPSv2}}$; we used grid search with a granularity of 0.1 to optimize the accuracy of image preference prediction. Since the scales of the scores are different, all of them were standardized based on the mean and standard deviation of the training set. We obtained the same results of $w_{\mathrm{IR}}=0.1$, $w_{\mathrm{Pick}}=0.6$, and $w_{\mathrm{HPSv2}}=0.3$ for all LOPO training sets.
5.2.4 AU4 Valence Score: Independent Preference Prediction for Selected Image Pairs
We first evaluate AU4 valence score on its ability to independently predict user preferences between pairs of images generated based on the same text prompt. Ground truth labels of preferences between images are derived from the image ranking at the end of each session, which resulted in user preference data of a total of 7068 image pairs from the 30 included participants.
Considering the randomness of AU activation and the imperfection of the AU estimation model, a small difference between two AU4 valence scores may not be informative for inferring the preference between the two associated images. Therefore, we apply a difference threshold $d$ for making predictions: only when the difference between the AU4 valence scores of two images is larger than or equal to the threshold $d$, do we predict that the user prefers the image with the higher AU4 valence scores; otherwise, we abstain from making a prediction.
There are 2 parameters to fit for the proposed algorithm: the exponential decay coefficient $k$ (as in Eq. 3) and the difference threshold $d$. We fit these two parameters with an LOPO procedure using grid search (from 0.1 to 2.0 with a granularity of 0.1 for $k$ and from 0.00 to 1.00 with a granularity of 0.02 for $d$) to optimize
@@ 5.2.4 AU4 Valence Score: Independent Preference Prediction for Selected Image Pairs
where $n_{\mathrm{correct}}$ and $n_{\mathrm{incorrect}}$ represents the numbers of correct predictions and incorrect predictions on preference of image pairs respectively.
@@
CLIP Score 60.58
Aesthetic Score 50.99
BLIP Score 52.98
ImageReward Score 63.11
PickScore 62.57
HPS V2 Score 62.93
Ensemble Baseline Score 67.09
AU4 Valence Score 74.86
Table 1: Accuracy of image preference prediction on the 15.6% of all data (1106 out of 7068 pairs) as selected by the AU4 valence model. (Image pairs with AU4 valence score difference below the trained difference threshold are not included.)
@@
The results are shown in Tab. 1. Based on the trained difference thresholds, the AU4 valence model makes predictions on 15.6% of the image preference data (1106 out of 7068 pairs). For the selected image pairs, our AU4 valence score has the best accuracy of 74.86%, outperforming the ensemble of the three pre-trained human preference scoring models by a margin of 7.77%.
Figure 4: Heatmap of the image preference prediction accuracy of AU4 valence score on the selected subset of all data. The y-axis represents different values of the exponential decay coefficient in the computation of AU4 valence score, and the x-axis represent different values of the difference threshold. As indicated by the color bars, for the color of each pixel, the hue represents the AU4 valence score prediction accuracy associated with the set of two parameters (exponential decay coefficient and difference threshold) while the saturation represents the proportion of selected data under the set of two parameters. The orange square marks the pixel representing the set of coefficients with optimized Eq. 5 (hue of this pixel is used in the color bar for proportion of selected data). The color bar for accuracy ranges only from 0.5 to 1.0 in our visualization, as no accuracies below 0.5 were obtained.
Our results have high stability. Firstly, the parameter values fit in the LOPO procedure were generally consistent. We obtained the same values of $k=0.4$ and $d=0.42$ for 28 out of the 30 total training sets; for the other two sets, we had $k=0.5$ and $k=0.46$. Secondly, Fig. 4 shows the accuracies of AU4 valence score under different sets of the two parameters (exponential decay coefficient and difference threshold), from which we can clearly see that the accuracy is smooth in these parameters and that high accuracy holds for a wide range of parameters.
@@
5.2.5 Integrate AU4 Valence Score with Pre-trained Scoring Models
@@
AU4 valence score has a great performance of 74.86% on the subset of the image pairs with distinguishable AU4 valence scores, but if we force it to make a prediction on every image pair in the dataset, its accuracy falls to 56.62%, lower than that of the CLIP score, ImageReward score, PickScore, HPS v2 score, and ensemble baseline score (see Tab. 2).
@@
However, besides making independent predictions, the AU4 valence score can also be integrated with pre-trained scoring models to enhance their consistency with human preferences, which will be applicable to all generated images with valid facial expression reaction data. We integrate the pre-trained scoring models with AU4 valence score with a simple linear combination:
$s_{m}+a_{m}s_{\mathrm{AU4}},$ (6)
@@
where $m$ represents the specific pre-trained scoring model, $s_{m}$ the score given by the model $m$, and $a_{m}$ the weight for integrating model $m$ and the AU4 valence score.
@@
We separately investigated the effect of integrating the AU4 valence score with the three scoring models trained on large human preference datasets, ImageReward, PickScore, and HPS v2, and their ensemble. Here, we only fit the weights $a_{\mathrm{IR}}$, $a_{\mathrm{Pick}}$, $a_{\mathrm{HPSv2}}$, and $a_{\mathrm{ens}}$ individually with an LOPO procedure using grid search (from 0.0 to 3.0 with a granularity of 0.1) to optimize the accuracy of image preference prediction; for the weights of the ensemble baseline score (Eq. 4) and the exponential decay coefficient for the AU4 valence score (Eq. 3), we directly used the values fit with a similar LOPO procedure in Secs. 5.2.3 and 5.2.4.
Model Ind. Acc. Acc. w/ AU4
@@
AU4 Valence Score 56.62 /
ImageReward Score 60.44 62.68
PickScore 63.36 65.31
HPS V2 Score 61.94 65.45
Ensemble Baseline Score 65.79 66.81
Table 2: Accuracy of image preference prediction on all data. The second column shows the accuracy of the models given in the first column when they make predictions independently while the third column shows the accuracy of the best few when they are integrated with the AU4 valence score.
The results are shown in Tab. 2. Integration with AU4 valence score improves the performance of all of the four baseline models. Specifically, by further integrating all three pre-trained human preference scoring models together with the AU4 valence score, we achieve the highest accuracy of 66.81%, outperforming the ensemble of the three models by a margin of 1.02%.
5.3 Aspects of Evaluations Reflected
We have just compared the performances of different scoring models. An interesting question is what aspects of the evaluation these models respectively reflect. Evaluations on image-text alignment and fidelity (image quality) are the two most essential aspects of text-to-image generation assessment researchers have been focusing on [52]. In our annotation survey, participants give a rating for each of them on a scale of 1 to 7 in addition to the overall rating of the image. To investigate their respective effects on the AU4 valence score, we apply a multiple linear regression from image-text alignment ratings and fidelity ratings to standardized scores from different models across all of the 30 participants included in Sec. 5.2. (AU4 valence scores are computed with a consistent exponential decay coefficient of 0.4; as mentioned in Sec. 5.2.4, the same parameter value of $k=0.4$ was obtained consistently in 28 out of the 30 LOPO training sets.) As shown in Tab. 3, the AU4 valence score has the most significant coefficient for fidelity ratings while the pre-trained scoring models (ImageReward, PickScore, HPS v2, and their ensemble) have more significant coefficient for image-text alignment ratings. This suggests that the aspect of evaluation AU4 primarily reflects is complementary to that of the pre-trained scoring models and explains why integrating the AU4 valence score with them consistently improves their performance.
Model Coef. for Alignment P-val. for Alignment Coef. for Fidelity P-val. for Fidelity
CLIP Score $-3.15\times 10^{-2}$ $5.77\times 10^{-2}$ $7.47\times 10^{-2}$ $2.33\times 10^{-5}$
Aesthetic Score $-8.71\times 10^{-3}$ $6.01\times 10^{-1}$ $1.85\times 10^{-2}$ $2.97\times 10^{-1}$
BLIP Score $-1.95\times 10^{-2}$ $2.41\times 10^{-1}$ $1.72\times 10^{-2}$ $3.31\times 10^{-1}$
ImageReward Score $2.53\times 10^{-1}$ $3.71\times 10^{-57}$ $-5.30\times 10^{-2}$ $1.19\times 10^{-3}$
PickScore $1.53\times 10^{-1}$ $9.06\times 10^{-22}$ $3.54\times 10^{-2}$ $3.53\times 10^{-2}$
HPS V2 Score $6.98\times 10^{-2}$ $2.31\times 10^{-5}$ $3.49\times 10^{-2}$ $4.64\times 10^{-2}$
Ensemble Baseline Score $1.38\times 10^{-1}$ $6.64\times 10^{-24}$ $2.64\times 10^{-2}$ $6.67\times 10^{-2}$
AU4 Valence Score $3.57\times 10^{-2}$ $2.89\times 10^{-2}$ $7.50\times 10^{-2}$ $1.64\times 10^{-5}$
Table 3: The results of the multiple linear regression from the image-text alignment ratings and fidelity ratings to standardized scores from different models. The five columns represent the name of the model, the coefficient for image-text alignment rating, the p-value for image-text alignment rating, the coefficient for fidelity rating, and the p-value for fidelity rating respectively.
6 Related Work
Facial Expression Recognition. Automatic facial expression recognition has advanced rapidly in recent years [41]. As emotional responses to image generation can be more complicated than single categories of the most widely-researched seven basic emotions [18, 17, 46], we decided to analyze the facial expression reactions directly in terms of muscle movements, as defined by AUs in FACS [19]. AU detection and estimation has attracted increasing interest [26, 61, 82, 81, 10, 42, 43, 69, 33, 34, 83, 45, 65, 66] and been used for facial emotion recognition [75, 77, 76, 74]. Specifically related to our usage of AU4 to infer user evaluations of generated images, researchers have shown that the activity of the corrugator supercilii (activated in AU4) is positively associated with amygdala and negatively associated with ventromedial prefrontal cortex activity [23] and increases with negative stimuli, such as negative images [5, 4, 37, 36].
@@ 7 Discussion and Conclusion
We propose that annotation of user preferences for text-to-image generation can be automated with analysis of user facial expression reaction to generated images, which can potentially increase the efficiency and scalability of human preference data collection. We present the FERGI dataset, comprising of video recordings of facial expression reaction to text-to-image generation, and show that multiple AUs are correlated with the participant’s evaluation of and emotional reaction to the generated image.
Specifically, AU4 activation is most consistently reflective of negative evaluation of the image, so we developed an AU4 valence score to represent the user’s evaluation of the image reflected by AU4 activation, which is useful in automatically annotating user preferences between pairs of images generated from the same input text prompt. Firstly, the AU4 valence score independently can select a subset of available data with large score difference and achieve a high accuracy in preference annotation of the selected data, substantially outperforming state-of-the-art pre-trained scoring models, including the ensemble of them. Therefore, it can potentially be used to selectively annotate large-scale user preference data to be used for fine-tuning scoring models that do not rely on data of facial expression reaction. Secondly, the AU4 valence score can be integrated with pre-trained scoring models to improve their overall accuracy in image preference prediction, which can be potentially helpful when the scoring models are used to label data of large scale human preferences for fine-tuning text-to-image generative models. Thirdly, AU4 may be useful for understanding individual user preferences to create personalized image generation models. Additionally, we show that the AU4 valence score is the most reflective of the evaluation of the fidelity of the generated images among all models, which is complementary to the pre-trained scoring models, which better reflect the evaluation of image-text alignment.
Finally, we have demonstrated the feasibility of automatically annotating user preferences for image generation from facial expression reaction in this work. The application of this method is not limited to text-to-image generation but may also be potentially applicable to other image generation tasks, such as image-to-image translation [53], image inpainting [20], and super-resolution [70].
@@ 8 Acknowledgments
A Supplementary Visuals
Figure S1: A flow chart summarizing the pipeline our paper proposes. The solid lines represent the procedures demonstrated in the paper while the dashed lines represent the procedures proposed but not demonstrated in the paper. Specifically, we developed a model and method for computing an AU4 valence score and integrating it with state-of-the-art text-to-image scoring models. We propose that the AU4 valence score could be used to fine-tune/improve these text-to-image scoring models and that the integrated ensemble score could be used to improve/fine-tune the text-to-image generative models.
Figure S2: An example of AU4 activation in response to a low-quality image generation. (a) The input text prompt. (b) The facial expression of the participant before seeing the generated image. (c) The generated image. (d) The facial expression of the participant after seeing the generated image (the frame with the highest estimated AU4 intensity). In the annotation survey, the participant gives an overall rating of 2, an image-text alignment rating of 2, and a fidelity rating of 1; she identified the image with the issues “output contains unwanted content that was not mentioned in the text prompt” and “existence of body problem”; she reported that she felt “disappointed” when seeing the image; she also ranked the image as the worst among the five images generated from the same prompt.
To better illustrate the idea of our paper, we present Fig. S1 showing a flow chart summarizing the pipeline the paper proposed and Fig. S2 showing an example of a strong AU4 activation in response to a low-quality image generation.
B Data Collection
@@
This section provides more details of the collection of the FERGI dataset not provided in Sec. 2. Figs. S3, S4, S5, S6, S7, S8 and S9 display screenshots of an example of one session in data collection.
B.1 Participants
@@
39 participants [^1] The user IDs are up to 040 as shown in Tab. S1 because the same participant used both 027 and 028 due to some technical issues. Though, no data was collected under the user ID 027. were recruited from the SONA system of University of California San Diego (UCSD). They underwent the informed consent process and received instruction for the study from the researcher via a videoconference and completed the study asynchronously afterwards on their personal computers. The videos of 3 participants failed to record and displayed only a logo of OBS Studio (the reason has not yet been fully understood by the researchers); 1 participant dropped out of study under the instruction of the researcher because the only available personal device was a tablet, which is not compatible with some procedures of the study; 1 participant did not complete any study sessions asynchronously at all after receiving the instruction in the videoconference. Among the other 34 participants with valid data and videos recorded, only 1 participant chose not to permit video recordings being shared with other researchers for research purposes; therefore, we decided to also exclude the data of this participant in our analysis for enhancing reproducibility of our study within the research community. We ended up with a dataset of 33 participants that is available upon request for research purposes.
B.2 Prompt Creation

### 21 Mar 2024 v2

@@ 6.2.3 Baseline Scores
To compare the performance of our AU valence score, we use the CLIP score [55], Aesthetic score [60], BLIP score [41], ImageReward score [75], PickScore [33], and HPS v2 score [73] as baselines.
@@
$s_{\mathrm{ens}}=w_{\mathrm{IR}}s_{\mathrm{IR}}+w_{\mathrm{Pick}}s_{\mathrm{Pick}}+w_{\mathrm{HPSv2}}s_{\mathrm{HPSv2}}$ (7)
@@
To be consistent with the evaluation of AUcomb valence score, a leave-one-participant-out (LOPO) procedure was applied in the fitting of the three weights $w_{\mathrm{IR}}$, $w_{\mathrm{Pick}}$, and $w_{\mathrm{HPSv2}}$; we used grid search with a granularity of 0.1 to optimize the accuracy of image preference prediction. Since the scales of the scores are different, they were each z-scored using their mean and standard deviation on the training set. We obtained the same results of $w_{\mathrm{IR}}=0.1$, $w_{\mathrm{Pick}}=0.6$, and $w_{\mathrm{HPSv2}}=0.3$ for all LOPO training sets.
6.2.4 AUcomb Valence Score: Independent Preference Prediction for Selected Image Pairs
We first evaluate the AUcomb valence score on its ability to independently predict user preferences between pairs of images generated by the same text prompt. Ground truth labels of preferences between images are derived from the image ranking at the end of each session, which resulted in user preference data on a total of 3534 image pairs from the 30 included participants.
Considering the randomness of AU activation and the imperfection of the AU estimation model, a small difference between two AUcomb valence scores may not be informative for inferring preference between the two associated images. Therefore, we apply a threshold $d$ for making predictions: only when the difference between the AUcomb valence scores of two images is larger than or equal to the threshold $d$, do we predict that the user prefers the image with the higher AUcomb valence score; otherwise, we abstain from making a prediction.
There are 3 parameters to fit for the proposed algorithm: the exponential decay coefficient $k$ (as in Eqs. 3 and 4), the weight for AU4 $w_{\mathrm{AU4}}$ (as in Eq. 5; $w_{\mathrm{AU12}}$ is then automatically set by Eq. 6) and the difference threshold $d$. We fit these three parameters with an LOPO procedure using grid search (from 0.1 to 2.0 with a granularity of 0.1 for $k$, from 0.0 to 1.0 with a granularity of 0.1 for $w_{\mathrm{AU4}}$, and from 0.00 to 1.00 with a granularity of 0.02 for $d$) to optimize
@@ 6.2.4 AUcomb Valence Score: Independent Preference Prediction for Selected Image Pairs
where $n_{\mathrm{correct}}$ and $n_{\mathrm{incorrect}}$ represents the number of correct predictions and incorrect predictions on image pair preference.
Table 1: Accuracy of image preference prediction on the 19.38% of all data (685 out of 3534 pairs) as selected by the AUcomb valence model. (Image pairs with AUcomb valence score difference below the trained difference threshold are not included.)
@@
CLIP Score 58.25
Aesthetic Score 51.39
BLIP Score 49.78
ImageReward Score 62.19
PickScore 61.31
HPS V2 Score 62.04
Ensemble Baseline Score 66.57
AUcomb Valence Score 72.70
@@
The results are shown in Tab. 1. Based on the trained difference thresholds, the AUcomb valence model makes predictions on 19.38% of the image preference data (685 out of 3534 pairs). For the selected image pairs, our AUcomb valence score has the best accuracy of 72.70%, outperforming the ensemble of the three pre-trained human preference scoring models by a margin of 6.13%.
Our results have high stability. The parameter values fit in the LOPO procedure were generally consistent. We obtained $k=0.5$, $w_{\mathrm{AU4}}=0.5$, $w_{\mathrm{AU12}}=0.5$, and $d=0.26$ for 20 out of the 30 total training sets.
@@
6.2.5 Integrate AUcomb Valence Score with Pre-trained Scoring Models
@@
AUcomb valence score has a great performance of 72.70% on the subset of the image pairs with distinguishable AUcomb valence scores, but if we force it to make a prediction on every image pair in the dataset, its accuracy falls to 58.23%, still significantly better than chance (two-sided binomial test: $p<5\times 10^{-22}$) but lower than that of the ImageReward score, PickScore, HPS v2 score, and ensemble baseline score (see Tab. 2).
@@
However, besides making independent predictions, the AUcomb valence score can also be integrated with pre-trained scoring models to enhance their consistency with human preferences, which will be applicable to all generated images with valid facial expression reaction data. We integrate the pre-trained scoring models with AUcomb valence score with a simple linear combination:
$s_{m}+a_{m}s_{\mathrm{AUcomb}},$ (10)
@@
where $m$ represents the specific pre-trained scoring model, $s_{m}$ the score from model $m$, and $a_{m}$ the weight for integrating $m$ and the AUcomb valence score.
@@
We separately investigated the effect of integrating the AUcomb valence score with each of the three scoring models trained on large human preference datasets (ImageReward, PickScore, and HPS v2), and their ensemble. Here, we only fit the weights $a_{\mathrm{IR}}$, $a_{\mathrm{Pick}}$, $a_{\mathrm{HPSv2}}$, and $a_{\mathrm{ens}}$ individually with an LOPO procedure using grid search (from 0.0 to 10.0 with a granularity of 0.1); for the weights of the ensemble baseline score (Eq. 7), exponential decay coefficient for the AU4 and AU12 valence scores (Eqs. 3 and 4), and the weighting between AU4 and AU12, we directly used the values fit previously.
Table 2: Accuracy of image preference prediction on all data. The second column shows the accuracy of the models given in the first column when they make predictions independently while the third column shows the accuracy of the best few when they are integrated with the AUcomb valence score.
Model Ind. Acc. Acc. w/ AUs
@@
AUcomb Valence Score 58.23 /
ImageReward Score 60.44 63.67
PickScore 63.36 65.42
HPS V2 Score 61.94 65.53
Ensemble Baseline Score 65.79 67.49
The results are shown in Tab. 2. Integration with AUcomb valence score improves the performance of all of the four baseline models. Specifically, by further integrating all three pre-trained human preference scoring models together with the AUcomb valence score, we achieve the highest accuracy of 67.49%, outperforming the ensemble of the three models by a margin of 1.70%.
@@ 7 Discussion and Conclusion
We propose that annotation of user preferences for text-to-image generation can be automated with analysis of user facial expression reaction to generated images, which can potentially increase the efficiency and scalability of human preference data collection. We present the FERGI dataset, consisting of video recordings of facial expression reaction to text-to-image generation, and show that multiple AUs are correlated with the participant’s evaluation of and emotional reaction to the generated image.
Specifically, we developed an AUcomb valence score to represent the user’s evaluation of the image reflected by AU4 and AU12 activations, which is useful in automatically annotating user preferences between pairs of images generated from the same input text prompt.
Firstly, the AUcomb valence score independently can select a subset of available data with large score difference and achieve a high accuracy in preference annotation of the selected data, substantially outperforming state-of-the-art pre-trained scoring models, including their ensemble. Therefore, it can potentially be used to selectively annotate large-scale user preference data used for fine-tuning scoring models. Note that automatic annotation will enable collection of vastly more human preference data from real users, so even if only a small portion are usable, its scale will still be much larger than manually annotated datasets.
Secondly, the AUcomb valence score can be integrated with pre-trained scoring models to improve their accuracy in image preference prediction, which can be potentially helpful when the scoring models are used to label human preferences for fine-tuning text-to-image generative models. Importantly, previous work has shown that even if the accuracy of the predictions is not very high, it can still be effective for fine-tuning of text-to-image models: Xu et al. [75] used the reward model ImageReward (trained from manually labeled human preferences), which only has an accuracy of 65.14% on their own dataset and 60.44% on our FERGI dataset, to fine-tune SD v1.4 and obtained a new model whose generated images were preferred by humans 58.79% of the time over those generated by the original SD v1.4. (see Tab. 4 and Fig. 6 of [75]).
Thirdly, our AUcomb valence score may be useful for understanding individual user preferences to create personalized image generation models.
In conclusion, we have demonstrated the feasibility of automatically annotating user preferences for image generation from facial expression reaction. The application of this method is not limited to text-to-image generation but may also be potentially applicable to other image generation tasks, such as image-to-image translation [53], image inpainting [21], and super-resolution [70].
@@ 8 Acknowledgments
A Data Collection
@@
This section provides more details of the collection of the FERGI dataset not provided in Sec. 3. Figs. S1, S2, S3, S4, S5, S6 and S7 display screenshots of an example of one session in data collection.
A.1 Participants
@@
39 participants [^1] The user IDs are up to 040 as shown in Tab. S1 because the same participant used both 027 and 028 due to some technical issues. Though, no data was collected under the user ID 027. were recruited from the SONA system of University of California San Diego (UCSD). They underwent the informed consent process and received instruction for the study from the researcher via a videoconference and completed the study asynchronously afterwards on their personal computers. The videos of 3 participants failed to record and displayed only a logo of OBS Studio (the reason has not yet been fully understood by the researchers); 1 participant dropped out of study under the instruction of the researcher because the only available personal device was a tablet, which is not compatible with some procedures of the study; 1 participant did not complete any study sessions asynchronously at all after receiving the instruction in the videoconference. Among the other 34 participants with valid data and videos recorded, only 1 participant chose not to permit video recordings being shared with other researchers for research purposes; therefore, we decided to also exclude the data of this participant in our analysis for enhancing reproducibility of our study within the research community. We ended up with a dataset of 33 participants that will be made available to researchers for research purposes.
A.2 Prompt Creation